OpenStudy (anonymous):

What is the general solution of tan(3θ)=cot(2θ)?

OpenStudy (anonymous):

\[\tan 3 \theta = \cot 2 \theta\]\[\tan 3 \theta = \tan (\frac{\pi}{2} - 2 \theta )\]\[3 \theta = \frac{\pi}{2} - 2 \theta + k \pi\]\[\theta = \frac{\pi}{10} + k \frac{\pi}{5}\]k is an integer